OpenStudy (anonymous):

3 over x minus the fraction 3 over x plus 1 plus the fraction 3 over x plus 2.

OpenStudy (anonymous):

= 3(x+1)(x+2)-3x(x+2)-3x(x+1) x(x+1)(x+2) = 3(x^2+3x+2-x^2-2x-x^2-x) x(x+1)(x+2) = 3(2-x^2) x(x+1)(x+2)

Parth (parthkohli):

Answer is 3 3/x

OpenStudy (anonymous):

how did u get that

Parth (parthkohli):

Okay, let me draw

Parth (parthkohli):